Sodium reacts with water to produce hydrogen gas and sodium hydroxide. How many grams of H2(g) are produced when 5.0 g Na reacts with 100.0 g of water? 


1
Expert's answer
2022-05-02T06:05:02-0400

"2Na+2H_2O \rightarrow 2NaOH + H_2"

M(Na)=23 g/mol

n(Na)=5/23=0.217 mol

"M(H_2O)=1x2+16=18 g\/mol"

"n(H_2O)=100\/18=5.556 mol"

As we can see from reaction, "[Na]:[H_2O]=1:1"

So Na is limited reagent. From reaction we know that "n(H_2)=n(Na)\/2=0.109 mol"

"M(H_2)=2x1=2 g\/mol"

"m(H_2)=M(H_2)xn(H_2)=2x0.109=0.217 g"